Tiger Golf Supplies has $19 million in earnings with 6 million shares outstanding. Its investment banker thinks the stock should trade at a P/E ratio of 22. If there is an underwriting spread of 9.4 percent.

a. What should the price to the public be? (Enter the answer in dollars not in millions. Round the intermediate calculations to two decimal places. Round the final answer to 2 decimal places.) Price $

b. What would the firm net? (Round the intermediate calculations to two decimal places. Round the final answer to 2 decimal places.) Net amount received
Answer #1

From the desk of DATE With 19 million in eamings & 6 million shaies, we fust compute the EPS. O EPS Camings / number of shape
Answer #2
19/6=3.167*22=69.67 I hope this helps
answered by: anonymous